# Nonparametric regression for cost-effectiveness analyses with observational data - a tutorial

### v1

@@ 4.2 BART
where each $f_{i}$ is a regression tree. Thereby the model combines the contributions of many small regression trees, each capturing different patterns in the data. This approach allows for modelling complex nonlinear relationships between covariates and outcome without making strong parametric assumptions. $200$ are frequently used, since this choice has been found to provide good performance while keeping the computational demands of the model fitting manageable. See Chipman et al. (2010); Tan and Roy (2019) for detailed treatments of the original BART model, and in particular the process by which the trees are learned from the data.
@@
The acronym suBART stands for seemingly unrelated BART, a name deliberately reminiscent of the SUR model we described above. SUR consists of two linear models which are coupled through correlated error terms. suBART keeps the latter, but replaces the linear models with additive regression trees. More precisely, the model is given by:
$\begin{pmatrix}c_{i}\\ q_{i}\end{pmatrix}$ $=\begin{pmatrix}\sum^{m}_{j=1}f_{j}(\mathbf{x_{i}})\\ \sum^{m}_{j=1}g_{j}(\mathbf{x_{i}})\end{pmatrix}+\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix},$ $\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix}\sim\text{N}\Biggl{(}\begin{pmatrix}0\\ 0\end{pmatrix},\boldsymbol{\Sigma}\Biggl{)},$ (2)
@@ 4.4 A note on Bayesian estimation and MCMC
Bayesian estimation has a long tradition in CEA, with some authors arguing it is the most natural framework to perform cost-effectiveness research under (Baio, 2018). The main reason is that cost-effectiveness researchers want to speak about the probability of cost-effectiveness which is nothing other than $\Pr(\lambda\Delta_{q}-\Delta_{c}>0)$. In other words, we want to make probabilistic statements about our parameters of interest, and this – strictly speaking — is only possible under the Bayesian paradigm.
@@
In addition, there are pragmatic considerations which motivate our use of Bayesian estimation: it enables the use of complex statistical models which would not be usable otherwise, and which perform well even with respect to non-Bayesian criteria. BART is a prime example of this, as shown through the results of the American Causal Inference Conference (ACIC) competition (Dorie et al., 2019).
@@
In practice, Bayesian estimation works by taking a large number of random draws from the posterior distribution. We then use these draws to approximate our quantities of interest, such as the aforementioned probability of cost-effectiveness or confidence intervals for the treatment effects. This approximation will be quite accurate if the number of draws is sufficiently large. One popular procedure to obtain these draws is is called Markov chain Monte Carlo (MCMC), and the collection of draws is called an MCMC sample. The use of MCMC samples is quite similar to that of bootstrap samples, even though the underlying theory is different. For example, to construct a 90% confidence interval, we can simply take the 5th and 95th percentiles of the MCMC sample. Also analogously to bootstrapping, there are no strict guidelines for what the aforementioned ”large number” of draws should be. The usual choice is a few thousand, and we will use $4000$ in the example below.
Details on the specific MCMC method for the suBART model may be found in Esser et al. (2024).
4.5 ps-suBART
@@
In the causal inference literature, one frequently encounters the so-called propensity score (PS), which is the probability that patient received the treatment, given their baseline characteristics. There PS It summarizes all the confounding variables into a single number. The inclusion of the PS as an additional covariate is very beneficial when estimating treatment effects through BART (Hahn et al., 2020; Dorie et al., 2019). In practice, we first estimate the PS for every patient in the data — for example, a BART model that predicts treatment assignment. We then add this PS to the covariates when fitting the BART model for the outcomes. This idea extends in the naturally to suBART, and the resulting ps-suBART model has been shown to perform well in the setting observational CEAs (Esser et al., 2024).
@@ 5 Guided example
In this section we guide the reader through the use of suBART. We use a simulated dataset, which is intended as an example and not meant to represent any specific real-world research endeavour. In the supplementary material [^6] See https://github.com/Jonas-Esser/nonparametric_CEA, we provide the full dataset, together will all of the following code. The dataset contains the treatment indicator (which indicates which treatment a patient is part of), the two outcome variables (costs and health), and multiple confounding variables. FOr illustration, we show part of the data in Table 1.
Table 1: Example data.
@@
In order for the following code to work correctly, it is important that the data are formatted properly. Unordered categorical variables with more than two levels (i.e., variables that represent categories without any natural order, such as blood type, region, or treatment center) should be declared as factors. In the simulated dataset, this applies to the trauma_type variable. All other variables should be declared as numbers. Furthermore, all missing observations should be properly formatted as NA values, not as text strings or numbers such as $-99$, which is sometimes done.
1data\$education <- as.factor(data\$education)
2data[data == -99] <- NA
Listing 1: Data formatting
@@ 5.1 Data preparation
We put all the covariates into a data frame called X, including the treatment indicator but not the outcomes c and q. The X object should contain only those variables which are supposed to enter into the regression model, and no others. Likewise, we put the outcomes c and q into their own data frame Y.
@@
Listing 2: Data preparation
@@ 5.2 Estimation of propensity scores
We now proceed to estimate the propensity scores (PS). The number of trees and MCMC iterations has been motivated earlier. We also use a so-called burn-in sample of size $1000$; the reasons for this need not concern us here and can be found in any introduction to MCMC.
@@
10
11X_ps <- X
12X_ps\$ps <- apply(pnorm(ps_fit\$y_hat_test), 1, mean)
Listing 3: Estimation of propensity scores
@@ 5.3 Further data preparation
As mentioned above, nonlinear models like BART have the property that the treatment effects estimates are not represented by a single model parameter, as is the case in the simple linear model without interaction terms. The treatment effects are instead computed based on predictions from the model, as follows: for each patient, we estimate the expected outcomes under both treatment options and calculate then the average difference between these expected outcomes. The procedure is commonly known as g-computation (Hernán and Robins, 2024).
We shall now lay out how this is done. The idea is to compute the expected outcomes for each patient under both treatment options, and then average the difference between these expected outcomes. More precisely, our strategy may be outlined as follows:
@@
    For each patient $i$ predict, according to the model:
    • $\hat{c_{i}}(0)$ and $\hat{q_{i}}(0)$, which are the expected costs and QALY of the patient, had they treatment 00, regardless of what treatment they actually received.
    • $\hat{c_{i}}(0)$ and $\hat{q_{i}}(0)$; defined analogously
@@
    Compute the incremental net benefit for any willingness-to-pay $\lambda$ of interest, through
@@
These steps will be performed for every MCMC draw. For the sake of step $3$, we create another data frame, which contains contains each patient twice: once with the treatment indicator set to 00, and once with it set to $1$. This data frame is called X_test in the code.
@@
Listing 4: Further data preparation
@@ 5.4 Fitting the suBART model
1suBART_ps_fit <- subart(x_train = X_ps,
@@
Listing 5: Fitting suBART
The resulting suBART_fit object contains a variety of objects; we shall only need the entry suBART_fit\$y_hat_test, which are the predictions from Step 5.3.
@@ 5.5 Obtaining the results
In this step we apply the calculations from Section 5.3 to all draws in the MCMC sample. We then use the thereby obtained draws of $\Delta_{c}$ and $\Delta_{q}$ to estimate $\Pr\left(\text{INB}_{\lambda}>0\right)$ for a specified range of $\lambda$ values.
1suBART_ps_fit\$ATE <- matrix(NA, nrow = 2000, ncol = 2)
2for (i in 1:2000){
3 suBART_ps_fit\$ATE[i,] <- c(
4 mean(suBART_ps_fit\$y_hat_test[X_test\$t == 1,1,i]) - mean(suBART_ps_fit\$y_hat_test[X_test\$t == 0,1,i]),
5 mean(suBART_ps_fit\$y_hat_test[X_test\$t == 1,2,i]) - mean(suBART_ps_fit\$y_hat_test[X_test\$t == 0,2,i])
6
@@
8}
9
10MCMC_sample <- data.frame(
11 Delta_c = suBART_ps_fit\$ATE[,1],
12 Delta_q = suBART_ps_fit\$ATE[,2]
13)
14MCMC_sample\$INB20 <- 20000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
15MCMC_sample\$INB50 <- 50000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
16
17lambda <- seq(0, 50000, length.out = 1000)
18p <- rep(NA, length(lambda))
19for (i in 1:length(lambda)){
@@ 12 Delta_q = suBART_ps_fit\$ATE[,2]
21}
22CEAC <- data.frame(lambda, p)
Listing 6: Postprocessing of MCMC output
@@ 5.6 Plotting the results
To finish the analysis, we plot the cost-effectiveness plane and the cost-effectiveness acceptability curve (Willan and Briggs, 2006), using the ggplot2 package (Wickham, 2016).
@@
3
4CEplane <- ggplot(MCMC_sample) +
@@
14
15
16CEAC <- ggplot(data = CEAC) +
@@
27
28grid.arrange(CEplane, CEAC, nrow = 1)
Listing 7: Plotting the results
Figure 3: Cost-effectiveness plane and CEAC.
@@ 6 Discussion
In this tutorial, we illustrated how to perform statistical estimation in cost-effectiveness analysis with observational data.
@@
We focused on the suBART model, both due to personal affinity with it, and because it has been shown to perform well in the aforementioned use case. We stress, however, that our main intention is not a presentation of this model, but rather to provide more applied cost-effectiveness researchers a more general framework for analysing observational data. suBART is merely an example, and there are alternatives which can be used in a similar way, and which we also expect to perform well. In principle, any non-Bayesian method for estimating $\Delta_{c}$ and $\Delta_{q}$ could be used, in conjunction with bootstrapping. The MCMC sample in Section 6 would then be replaced by the bootstrap sample, and the rest of the analysis proceeds in the same way. Such an analysis is no longer Bayesian, and hence the interpretation of the results changes: strictly speaking, we can no longer speak of the ”probability of cost-effectiveness”, and must instead argue in terms of frequentist p-values. See Löthgren and Zethraeus (2000) for details. Also, as mentioned earlier, such approaches may be infeasible in practice, due to computational constraints. Suitable choices for the estimator could include, for example, the Causal Forest of Wager and Athey (2018) or Double Machine Learning (Knaus, 2022).
We now discuss some limitations of our presented approach (which are however shared by the aforementioned alternatives) as well as possible remedies.
@@
  • suBART can only deal with missing observations in the outcome variables, not in the covariates. The technical reason for this is that the model is formulated conditional on the covariates, and hence makes no assumptions about the distribution of the covariates. In principle this is a good thing, because the analysis is therefore not dependent on such assumptions, but it also means that there is no way to impute missing covariates. Conceptually it seems most appealing to directly impute the missing covariates using the BART machinery, as proposed by Xu et al. (2016). While applying this idea to suBART is conceptually straightforward, a computational implementation is currently not available. As a pragmatic alternative, we suggest combining suBART with multiple imputation in the same way as is commonly done when performing CEAs with trial data. The procedure can be sketched as follows: we create multiple complete dataset using standard imputation techniques and compute, for each of these datasets, the point estimates and covariance matrix of $\Delta_{c}$ and $\Delta_{q}$ (see Listing 6). These are then pooled using Rubin’s rules, and the pooled estimates may be used to construct the CEAC. In this context we refer to Harel and Zhou (2007) for a general review of multiple imputation, and Ben et al. (2023) for specific guidance on imputation in trial-based CEAs.
  • We focussed on the basic setting in observational studies, where the treatment for each patient is assigned once and then not changed anymore. Most cost-effectiveness studies in the literature are of this kind, but there is also increasing interest in performing CEAs with so called time-varying treatments [add reference]. In such analyses, the patients may switch from one treatment to another during follow-up. The probability of such switches happening may depend on confounders, and these confounders may themselves change over time - this is accordingly called time-varying confounding. Statistical adjustment for such confounding is quite complicated; Hernán and Robins (2024) devote about a third of their book to it. One of their proposed methods, called the g-formula for time-varying treatments (Section 21.1) could in principle be combined with BART, and some first efforts in this direction have been made by Wang et al. (2024). Extending their work to the CEA setting seems possible in principle, but far beyond the scope of the present paper.
{acks}
[Acknowledgments] No financial support for this study was provided.

### v2

@@ 4.2 BART
where each $f_{i}$ is a regression tree; each of which captures different nonlinear effects of the covariates on the outcome. By constraining the trees to be small, BART effectively controls overfitting, which in turn improves generalization. This framework allows for modelling complex nonlinear relationships between covariates and outcome without relying on strong parametric assumptions. The total number of trees $m=200$ is frequently used, since this choice has been found to provide good performance while keeping the computational demands of the model fitting manageable. See Chipman et al. (2010); Tan and Roy (2019) for detailed treatments of the original BART model, and particularly the process by which the trees are learned from the data.
@@
The acronym suBART stands for seemingly unrelated BART, a name deliberately reminiscent of the SUR model we described in Section 3. SUR consists of two linear models which are coupled through correlated error terms. suBART keeps the latter, but replaces the linear models with additive regression trees. More precisely, the model is given by:
$\begin{pmatrix}c_{i}\\ q_{i}\end{pmatrix}$ $=\begin{pmatrix}\sum^{m}_{j=1}f_{j}(\mathbf{x_{i}})\\ \sum^{m}_{j=1}g_{j}(\mathbf{x_{i}})\end{pmatrix}+\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix},$ $\begin{pmatrix}\varepsilon_{i,1}\\ \varepsilon_{i,2}\end{pmatrix}\sim\text{N}\Biggl{(}\begin{pmatrix}0\\ 0\end{pmatrix},\boldsymbol{\Sigma}\Biggl{)},$
@@ 4.4 A note on Bayesian estimation and MCMC
Bayesian estimation has a long tradition in CEA, with some authors arguing that it is the most natural framework to perform cost-effectiveness research under (Baio, 2018). The main reason is that cost-effectiveness researchers are concerned with the probability of cost-effectiveness which is nothing other than $\Pr(\lambda\Delta_{q}-\Delta_{c}>0)$. In other words, we want to make probabilistic statements about our parameters of interest, while explicitly acknowledging uncertainty about their values, and this – strictly speaking — is only possible under the Bayesian paradigm.
@@
In addition, there are pragmatic considerations that motivate our use of Bayesian estimation. It allows for the use of complex statistical models that are computationally difficult to estimate using alternative approaches. Morever, Bayesian estimation often performs well even when assessed by frequentist or other non-Bayesian criteria. BART is a prime example, as shown through the results of the American Causal Inference Conference (ACIC) competition (Dorie et al., 2019).
@@
In practice, Bayesian estimation works by taking a large number of random draws from the posterior distribution. We then use these draws to approximate our quantities of interest, such as the aforementioned probability of cost-effectiveness or confidence intervals for the treatment effects. This approximation will be quite accurate if the number of draws is sufficiently large. One popular procedure to obtain these draws is is called Markov chain Monte Carlo (MCMC), and the collection of draws is called an MCMC sample. While superficially similar to bootstrapping, which is a commonly used tool in CEA, the underlying theory of MCMC is quite different. While bootstrap samples are resampled datasets used to approximate sampling variability in a frequentist framework, MCMC samples represent draws from the posterior distribution and quantify uncertainty in the parameters conditional on the observed data. In practice, however, the use of MCMC is quite similar to that of bootstrapping. For example, to construct a 90% credible interval [^6] The Bayesian analogue to the frequentist confidence interval., we can simply take the 5th and 95th percentiles of the MCMC sample. Also analogously to bootstrapping, there are no strict guidelines for what the aforementioned ”large number” of draws should be. The usual choice is a few thousand, and we will use 400040004000 in the example below. For more information on MCMC, we refer to Greenberg (2012), Chapter 7. Details on the specific MCMC method for the suBART model may be found in Esser et al. (2024).
4.5 Propensity scores
@@
In the causal inference literature, one frequently encounters the so-called propensity score (PS), which is the probability that a patient received the treatment, given their baseline characteristics. The PS summarizes all the confounding variables into a single number, and its inclusion as an additional covariate helps control for confounding more effectively (Hahn et al., 2020; Dorie et al., 2019). We therefore begin by estimating the PS for each patient in the data, using a BART model to predict treatment assignment based on baseline characteristics. This estimated PS is then included as an additional covariate when fitting the suBART model for the outcomes. This approach has been shown to perform well—and better than a suBART model without PS adjustment—in the context of observational CEAs (Esser et al., 2024).
@@ 5 Guided example
In this section we guide the reader through the use of suBART. We use a simulated dataset, which is designed as an example and not meant to represent any specific real-world research endeavour. In the supplementary material [^7] See https://github.com/Jonas-Esser/nonparametric_CEA, we provide the full dataset, together with all of the following code. The dataset contains the treatment indicator t (which indicates which treatment a patient received), the two outcome variables (costs and health — represented by c and q, respectively) and the confounding variables age, sex, and education. For illustration, we display a subset of the data in Table 1.
Table 1: Subset of the simulated dataset used in the guided example. Variables include treatment ($\mathbf{t}$), outcomes ($\mathbf{c}$, $\mathbf{q}$), and covariates (age,sex and education).
@@
5.1 Data formatting
In order for the suBART code to work correctly, it is important that the data are formatted properly. Unordered categorical variables with more than two levels (i.e., variables that represent categories without any natural order, such as blood type, region, or treatment center) should be declared as factors. In the dataset, this applies to the education variable. All other variables should be declared as numbers. Furthermore, all missing observations should be properly formatted as NA values, not as text strings or numbers such as 99-99- 99, which is occasionally observed in practice.
1rm(list=ls())
2set.seed(42)
3
4data_url <- "https://raw.githubusercontent.com/Jonas-Esser/nonparametric_CEA/refs/heads/main/data_example_missingness.csv"
5data <- read.csv(data_url)
6
7data\$education <- as.factor(data\$education)
8data[data == -99] <- NA
Listing 1: Loading the data, handling the unordered categorical variables and handle missing data in the outcomes, respectively.
@@ 5.2 Data preparation
We put all covariates into a data frame called X, including the treatment indicator but not the outcomes c and q. The X object should contain only those variables which are supposed to enter into the regression model, and no others. Likewise, we put the outcomes c and q into their own data frame Y.
@@
Listing 2: Construction of covariate data frame X and outcome data frame Y by selecting relevant covariates and response variables
@@ 5.3 Estimation of propensity scores
We proceed to estimate the PS using a BART model, where treatment assignment is modeled as a function of the previously defined confounders. The number of trees and MCMC iterations was specified earlier. In line with standard practice, we discard the first 1000 posterior draws (usually referred to as burn-in samples) following standard practice in the MCMC literature. This means that although we set the total number of MCMC iterations (n_mcmc) to 500050005000 in the code, only the remaining 400040004000 samples after burn-in are used for inference.
@@
9
10X_ps <- X
11X_ps\$ps <- apply(pnorm(ps_fit\$y_hat), 1, mean)
Listing 3: Estimation of PS by fitting a BART model using the subart package. When the outcome is univariate, subart reduces to fitting a standard BART model.
@@ 5.4 Further data preparation
As mentioned above, in nonlinear models such as BART, treatment effects estimates are not represented by a single model parameter, unlike with the simple linear model without interaction terms. Instead, they are derived from the model predictions as follows: for each patient, we estimate the expected outcomes under both treatment options and then calculate the average difference between these expected outcomes. The procedure is commonly known as g-computation (Hernán and Robins, 2024).
In algorithmic form, our strategy may be outlined as follows:
@@
    For each patient $i$, predict according to the model:
    • $\hat{c}_{i}(0)$ and $\hat{q}_{i}(0)$, which are the expected costs and QALYs of the patient, had they received treatment 0, regardless of the treatment they actually received.
    • $\hat{c_{i}}(1)$ and $\hat{q_{i}}(1)$; defined analogously.
@@
    Using the obtained $\Delta_{c}$ and $\Delta_{q}$, compute the incremental net benefit for any willingness-to-pay $\lambda$ of interest, through
@@
These steps will be performed for every MCMC draw. For the sake of step 333, we create another data frame, which contains each patient twice: once with the treatment indicator set to 0, and once with it set to 111. This data frame is called X_test in the code.
@@
Listing 4: Constructing the test dataset with treatment indicators to estimate potential outcomes under both treatment and control conditions.
@@ 5.5 Fitting the suBART model
1subart_fit <- subart(x_train = X_ps,
@@
Listing 5: Model fitting using suBART.
The resulting suBART_fit object contains a variety of objects; we only need the entry suBART_fit\$y_hat_test, which corresponds to the the predictions from Step 5.4.
@@ 5.6 Obtaining the results
In this step, we apply the calculations from Section 5.4 to all draws in the MCMC sample. We use the resulting draws of $\Delta_{c}$ and $\Delta_{q}$ to estimate $\Pr\left(\text{INB}_{\lambda}>0\right)$ for a specified range of $\lambda$ values, which will allow us to compute and plot the cost-effectiveness acceptability curve (CEAC), as described by Willan and Briggs (2006).
1n_post <- subart_fit\$mcmc\$n_mcmc - subart_fit\$mcmc\$n_burn
2subart_fit\$ATE <- matrix(NA, nrow = n_post, ncol = 2)
3
4# Calculating Step 5.4.3;
5for (i in 1:n_post){
6 subart_fit\$ATE[i,] <- c(
7 mean(subart_fit\$y_hat_test[X_test\$t == 1,1,i]) - mean(subart_fit\$y_hat_test[X_test\$t == 0,1,i]),
8 mean(subart_fit\$y_hat_test[X_test\$t == 1,2,i]) - mean(subart_fit\$y_hat_test[X_test\$t == 0,2,i])
9
@@
11}
12
13# Storing posterior MCMC sample
14MCMC_sample <- data.frame(
15 Delta_c = subart_fit\$ATE[,1],
16 Delta_q = subart_fit\$ATE[,2]
17)
18MCMC_sample\$INB20 <- 20000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
19MCMC_sample\$INB50 <- 50000 * MCMC_sample\$Delta_q - MCMC_sample\$Delta_c
20
21# Estimating probabilities for a specified range of lambda
22lambda <- seq(0, 50000, length.out = 1000)
23p <- rep(NA, length(lambda))
24for (i in 1:length(lambda)){
@@ 16 Delta_q = subart_fit\$ATE[,2]
26}
27CEAC <- data.frame(lambda, p)
Listing 6: Postprocessing of MCMC sample to compute average treatment effects, construct incremental $\text{INB}_{\lambda}$ at multiple willingness-to-pay thresholds, and calculate probabilities to build the CEAC
@@ 5.7 Plotting the results
Finally, we plot the cost-effectiveness plane and CEAC, using the ggplot2 package (Wickham, 2016).
@@
3CEplane <- ggplot(MCMC_sample) +
@@
13
14CEAC <- ggplot(data = CEAC) +
@@
25
26gridExtra::grid.arrange(CEplane, CEAC, nrow = 1)
Listing 7: Plotting the cost-effectiveness plane and CEAC based on the posterior MCMC sample.
Figure 3: Cost-effectiveness plane (left) and CEAC (right).
@@ 6 Discussion
In this tutorial, we demonstrated how to perform statistical estimation in CEAs with observational data. Our approach avoids strong parametric assumptions on the relationship between treatment and baseline variables with the expected costs and health state of a patient, which is very important when data is not obtained from a randomized trial. While we do not explicitly address the common issue of non-normal error terms, we argued that our approach is likely robust to violations of the normality assumption.
@@
We used the suBART model in our presentation, both because of its strong empirical performance in this context and its methodological advantages. We stress, however, that our primary goal was not to present this model, but rather to provide a general framework for cost-effectiveness researchers analyzing observational data. suBART is merely an example, and there are alternatives which can be used in a similar way, and which we also expect to perform well. In principle, any non-Bayesian method for estimating $\Delta_{c}$ and $\Delta_{q}$ can be used — for example, the Causal Forest of Wager and Athey (2018) or Double Machine Learning (Knaus, 2022). The MCMC sample referenced in Section 6 would then be replaced by a bootstrap sample, and the rest of the analysis be carried out analogously. However, such an analysis would be no longer Bayesian, and hence the interpretation of the results changes: strictly speaking, we could no longer speak of the ”probability of cost-effectiveness”, and must instead argue in terms of frequentist p-values; see Löthgren and Zethraeus (2000). Additionally, such approaches may be infeasible in practice due to computational constraints.
We now discuss some limitations of our presented approach, relating to missing data, and outline possible solutions to address them.
@@
  • suBART can only deal with missing observations in the outcome variables, not in the covariates. The technical reason for this is that the model is formulated conditional on the covariates, and hence makes no assumptions about their distribution. In principle this is a positive aspect, as the analysis does then not rely on such assumptions, but it also implies that there is no way to impute missing covariates. Conceptually it seems most appealing to directly impute the missing covariates using the BART machinery, as proposed by Xu et al. (2016). While applying this idea to suBART is conceptually straightforward, a computational implementation is currently not available. As a pragmatic alternative, we suggest combining suBART with multiple imputation in the same way as is commonly done when performing CEAs with trial data. The procedure can be sketched as follows: we create multiple complete datasets using standard imputation techniques and compute, for each of these datasets, the point estimates and covariance matrix of $\Delta_{c}$ and $\Delta_{q}$ (see Listing 6). These are then pooled using Rubin’s rules (Rubin, 2018), and the pooled estimates may be used to construct the CEAC. In this context, we refer to Harel and Zhou (2007) for a general review of multiple imputation, and to Ben et al. (2023) for specific guidance on multiple imputation in trial-based CEAs.
  • In our proposed framework, each patient’s cost and health outcomes are either fully observed or entirely missing. This excludes the common situation of censored observations, where data are partially missing. For example, patients may drop out of the study earlier than intended, and the probability of this dropout occurring may be related to the patients baseline characteristics as well as their health state and associated costs. Chen and Hoch (2022) proposed a framework to jointly address censoring and covariate adjustment in observational studies. However, their approach remains limited by its reliance on parametric regression models which must be specified correctly by the analyst. Their framework could in principle be combined with nonparametric methods for dealing with censoring — for example, the BART-based method by Sparapani et al. (2016). As in the previous bullet point, a computational implementation of such a method is currently lacking.
Despite these limitations, the presented framework is sufficient for use in many CEAs with observational data and produces more credible results than commonly used parametric models. We thus hope that readers will find it useful in their research.
7 Declarations
7.1 Funding
No funding was received for the completion of this paper.
7.2 Conflicts of interest
All authors have no conflicts of interest to declare.
7.3 Availability of data and materials
All data and code used are available under https://github.com/Jonas-Esser/nonparametric_CEA.
7.4 Ethics approval
Not applicable.
7.5 Author contributions
The initial idea for the paper was proposed by Jonas Esser; the scope and design were then worked out in cooperation with Johanna van Dongen and Judith Bosmans. The text was written primarily by Jonas Esser, while incorporating suggestions from all other authors. The R code, figures and computer experiments were created by Jonas Esser and Mateus Maia.